The shortest side of a right triangle measures 13m. The lengths of the other two sides are consecutive integers. Find the lengths of the other two sides
Found 2 solutions by Fombitz, Boreal:
Answer by Fombitz(32388)
Answer by Boreal(15235)

short side is 13
other two sides are x and x+1, and the longest is the hypotenuse.
(x+1)^2=x^2+13^2
x^2+2x+1=x^2+169
x^2 cancel
2x+1=169
2x=168
x=84 m
hypotenuse is 85 m
84^2 is 7056
85^2=7225
7056+169=7225, check